### v3

# Schema-Driven Information Extraction from Heterogeneous Tables

Abstract
In this paper, we explore the question of whether large language models can support cost-efficient information extraction from tables. We introduce schema-driven information extraction, a new task that transforms tabular data into structured records following a human-authored schema. To assess various LLM’s capabilities on this task, we present a benchmark comprised of tables from four diverse domains: machine learning papers, chemistry literature, material science journals, and webpages. We use this collection of annotated tables to evaluate the ability of open-source and API-based language models to extract information from tables covering diverse domains and data formats. Our experiments demonstrate that surprisingly competitive performance can be achieved without requiring task-specific pipelines or labels, achieving F1 scores ranging from 74.2 to 96.1, while maintaining cost efficiency. Moreover, through detailed ablation studies and analyses, we investigate the factors contributing to model success and validate the practicality of distilling compact models to reduce API reliance. [^1] Code and data are available at our GitHub repository.
Schema-Driven Information Extraction from Heterogeneous Tables
[…]
1 Introduction
Figure 1: Overview of Schema-Driven Information Extraction. The input includes two elements: the source code of a table and a human-authored extraction schema, outlining the target attributes and their data types. The output consists of a sequence of JSON records that conform to the extraction schema.
Vast quantities of data are locked away in tables found in scientific literature, webpages, and more. These tables are primarily designed for visual presentation, and the underlying data is typically not available in any structured format, such as a relational or graph database. Some table collections have simple or uniform structures Cafarella et al. (2008), making them easy to convert to relational data, for example, Wikipedia tables Lebret et al. (2016); Iyyer et al. (2017), however a lot of information is stored in tables with complex and varied layouts, such as tables of experimental data presented in scientific literature.
Prior work on extracting structured data from tables has focused on developing custom pipelines for each new table format or domain, for example extracting machine learning leaderboards from  result tables Kardas et al. (2020). Importantly, the development of these specialized pipelines necessitates domain-specific labeled data, which not only incurs a significant cost in collection for every new extraction task but also constrains their applicability outside the originating domain.
In this paper, we demonstrate that large language models can enable accurate domain-independent extraction of data from heterogeneous tables. To show this, we present a new formulation of the table extraction problem, which we refer to as Schema-Driven Information Extraction. In Schema-Driven IE, the only human supervision provided is a schema that describes the data model, including the target attributes and their data types, formulated in a JSON format. [^2] JSON is chosen as the output format for two main reasons: 1) its widespread use ensures a significant representation in the LLM’s pre-training corpus, which is crucial for optimizing model performance; and 2) its simplicity in parsing and processing, especially its support for one-line output, makes it advantageous for outputs spanning multiple cells, offering a clear benefit over indent-based formats like YAML. Given an extraction schema, and a table as input, the model then outputs a sequence of JSON objects, each of which describes a single cell in the table and adheres to the user-provided schema. For example, as demonstrated in Figure 1, a domain expert outlines the attributes of interest related to experimental result cells in a machine learning table, and the model extracts JSON objects following this schema.
To evaluate the ability of LLMs to perform Schema-Driven IE, we introduce a new benchmark consisting of table extraction datasets in four diverse domains: machine learning papers, chemistry literature, material science journals, and webpages - each of which has a different data format (, XML, CSV, and HTML, respectively). We curate and annotate data for the first two domains, while adapting existing datasets for the latter two.
We then use this benchmark to analyze the performance of open-source and proprietary LLMs. We find that proprietary models perform well across diverse domains and data formats. For example, GPT-4 OpenAI (2023) and code-davinci Chen et al. (2021), are capable of accurate table extraction (ranging from 74.2 to 96.1 F1), given only a relevant data schema as supervision. This performance is comparable to fully supervised models, which operate at an F1 range of about 64.1 to 96.1. We also present a number of analyses on various factors that are key to achieving good performance while minimizing inference costs, including retrieving text from outside the table, in addition to an iterative error recovery strategy. Moreover, we demonstrate the utility of Schema-Driven IE by evaluating performance on the downstream task of leaderboard extraction from machine learning papers (Kardas et al., 2020).
While open-source models have yet to match the performance of their proprietary counterparts, our analysis reveals that recent models like CodeLlama-instruct-13B (Rozière et al., 2023) show significant progress, e.g., on ML tables, its performance is comparable to that of GPT-3.5-turbo.
Furthermore, we conduct comprehensive ablation studies and analyses to explore model performance nuances, and demonstrate the feasibility of distilling efficient table extraction models without compromising performance. By introducing a new benchmark and a baseline for Schema-Driven IE, our goal is to encourage the creation of future open-source models capable of executing this task independently of proprietary APIs.
Figure 2: Left: Prompt formulation of our proposed method InstrucTE. Right: Illustration of our error-recovery strategy, which ensures the model compliance of the instructed cell traversal order and reduces inference costs.
2 Schema-Driven Information Extraction
We present a new task for extracting structured records from tables. As shown in Figure 1, the task input contains two elements: 1) a table with numerous cells, optionally supplemented with contextual text, e.g., retrieved paragraphs from the same document; and 2) an extraction schema that outlines target attributes and their data types for various record types (implemented as JSON templates). Given the input, the model generates a sequence of JSON objects, where each object corresponds to a cell in the table and contains key-value pairs for the pre-defined attributes of a specific record type.
Consider a table in an ML paper that displays various models’ results. Our proposed task enables the extraction of result records from each cell in the table. These records include relevant attributes such as the evaluation metric, task, etc, which are structured in corresponding JSON objects and could facilitate meta-analysis of experiments or support research on reproducibility.
To demonstrate the feasibility of Schema-Driven IE on tables, we introduce InstrucTE, a method to extract structured records from a broad range of semi-structured data, using only task-specific instructions. InstrucTE uses a template-based approach to information extraction (Chambers and Jurafsky, 2011; Chen et al., 2023), where the extraction schema is represented as a series of JSON templates. The underlying LLM is instructed to select the appropriate template and populate it with extracted values for each cell in an input table, following a specified cell traversal order. As illustrated in Figure 2 (left), the prompt used by InstrucTE consists of four key components: an input table (optionally) supplemented with contextual text, an extraction schema, task-specific instructions, and an initial record for starting the process.
Despite explicit instructions, we found that models often fail to generate JSON records for all the cells in a single inference pass. Instead, models deviate from the instructed cell traversal order, leading to partial extraction of the input table’s cells. To mitigate this, we use an iterative error recovery strategy. As shown on the right side of Figure 2, we detect deviations from the instructed left-right, top-down order by comparing predicted cell values with those from a rule-based cell detector. Then, we truncate the LLM’s output to the point of deviation, and re-prompt the model with the truncated sequence, adding the value of the next target cell. This strategy guides the model to adhere to the instructed order, and continues iteratively until all records are generated. In Section 4.4, we show that this approach achieves comparable performance to cell-by-cell prompting, while significantly reducing inference costs. For more details on InstrucTE, including prompt formulation and cell detectors, please refer to Appendix A.
3 The Schema-to-Json Benchmark
We now present the details of our benchmark, Schema-to-Json, which is designed to assess the capabilities of LLMs to extract data from tables, adhering to a predefined schema. This benchmark contains tables from four diverse domains: machine learning papers, chemistry literature, material science journals, and webpages. Each domain features a unique textual format, namely, , XML, CSV, and HTML, requiring models to generalize across domains and formats. For ML tables, we add relevant paragraphs from the same documents to provide additional context, testing the models’ capacity to jointly understand tabular and textual data. We manually annotate datasets for the first two domains and adapt pre-existing datasets into our unified format for the latter two. Statistics of the four datasets are summarized in Table 1.
arXiv Machine Learning Tables
We create a manually annotated dataset focused on tables from arXiv ML papers, emphasizing numeric cells that are classified into four categories: Experimental Results, Hyper-parameters, Data Statistics, or Other. Extraction attributes are pre-defined for the first three categories; for instance, Result records incorporate textual attributes such as evaluation metric (e.g., F1) and dataset (e.g., SQuAD), as shown in Figure 1. We collect papers from three subfields: Machine Learning, Computer Vision, and Natural Language Processing, and randomly select five tables from each paper (including those in appendices) for budget reasons. [^3] We limit the paper selection to those published between Oct. and Nov. 2022 to avoid contamination with GPT-4 (0613) and GPT-3.5, which were trained on data until Sep. 2021. We employ computer scientists with ML backgrounds for annotation, and evaluate inter-annotator agreement (IAA) score by calculating Table-F1 (a metric detailed in Section 4.1) on double-annotated tables, treating one set of annotations as gold labels and the other as predictions. This method yields a Table-F1 score of 96.6 when applying thresholded token-level F1 for attribute matching. For additional information on the ML tables, including predefined attributes and the annotation process, please refer to Appendix B.
PubMed Chemistry Tables
We also annotate a new dataset of PubMed tables describing the physical properties of chemical compounds. The automated extraction of physical properties from such tables could provide substantial real-world benefits, for example collecting much-needed data for training ML models that can support inverse molecular design (Kim et al., 2018) and thus accelerating the drug design process (Fields, 2019; Stokes et al., 2020). Here, we focus on cells concerning five important physical properties identified by chemists: $\text{IC}_{50}$, $\text{EC}_{50}$, $\text{GI}_{50}$, $\text{CC}_{50}$, and MIC. [^4] https://www.sciencedirect.com/topics/pharmacology-toxicology-and-pharmaceutical-science/ic50 Three common attributes are manually extracted from tables for all properties: unit, treatment (experimental compound), and target (measured biological entity, e.g., a gene expression). Similar to the ML tables, domain experts annotate JSON records for relevant cells, and Table-F1 calculated on double-annotated tables is used as the IAA score. A Table-F1 score of 91.0 is achieved when applying thresholded token-level F1 for attribute matching, underscoring the reliability of the dataset.
DisCoMat (Gupta et al., 2022)
We incorporate DisCoMat, an existing dataset focusing on glass composition tables from Elsevier material science journals. The task is to extract tuples comprising (material, constituent, percentage, unit) from given tables. We adapt DisCoMat to fit our Schema-Driven IE framework by grounding the percentage element to numeric cells in the table and considering the other elements as attributes. The model is tasked to identify numeric cells representing constituent percentages and predict the associated three attributes. We refer readers to Gupta et al. (2022) for more details of DisCoMat. [^5] In the released corpus, tables are represented as matrices; we, therefore, transform them into CSV tables (using the pipe symbol ”—” as the delimiter) prior to feeding them into LLMs.
SWDE (Hao et al., 2011)
Finally, we add SWDE (Structured Web Data Extraction) as a fourth dataset, aimed at extracting pre-defined attributes from HTML webpages. This dataset comprises roughly 124K pages gathered from eight distinct verticals, such as Auto, Book, and Movie. Each vertical includes ten unique websites and is associated with a set of 3 to 5 target attributes. For instance, the Movie vertical seeks to extract attributes such as title, director, and genre.
ML Chemistry DisCoMat SWDE
(ours) (ours) (2022) (2011)
Textual format XML CSV HTML
# cell types 4 6 2 8
# attr. types 11 4 4 32
# papers (web.) 25 16 2,536 80
# tables (pages) 122 26 5,883 1,600
# anno. records 3,792 1,498 58,481 1,600
# records / table 31.1 57.6 9.9 1
Table 1: Dataset statistics of four datasets in our Schema-to-Json benchmark.
4 Experiments
Figure 3: Capability of various LLMs to perform Schema-Driven IE, measured using the Schema-to-Json benchmark. We employ Table-F1 for our two newly annotated datasets and provide a measure of human performance. For DisCoMat (Gupta et al., 2022) and SWDE (Hao et al., 2011), we adhere to their original evaluation metrics, i.e., Tuple-F1 and Page-F1 respectively, to support comparisons with established methods. In SWDE experiments, k represents the number of trained websites from each vertical. Due to API cost constraints, *InstrucTE’s results are computed on a 1,600 webpage sample, with bootstrap confidence intervals calculated to validate the reliability of these performance estimates (margin of error for 95% confidence interval with 1000 samples is 0.00995.)
We evaluate the capability of various LLMs to perform schema-driven information extraction, in addition to full fine-tuning using our benchmark. For ML and chemistry tables, we use a subset of 10 and 7 randomly sampled papers separately for model development, which facilitates the training of supervised models, thereby enabling comparison with a schema-driven approach. For the two pre-existing datasets, we follow the data splits used in the original experiments.
4.1 Evaluation Metrics
We introduce Table-F1, a new reference-based evaluation metric gauging attribute prediction performance within a table. Table-F1 represents the harmonic mean of precision and recall, with precision being the ratio of correctly predicted attributes to total predicted attributes. At the attribute level, we consider two metrics: token-level F1 and exact match (EM). For token-level F1, a prediction is deemed correct if the score exceeds a specific threshold, which is determined by maximizing the alignment between model predictions and human judgments on the dev set (see Appendix C). An example of Table-F1 calculation is shown in Figure 7 in the appendix. We report macro-averaged Table-F1 given the wide variance in table sizes.
For DisCoMat and SWDE, we use the metrics specified in the original papers to support comparisons with prior work. In the case of DisCoMat, we report Tuple-F1 (Gupta et al., 2022), where a predicted 4-element tuple is considered correct only if it exactly matches the gold tuple. For SWDE, we report Page-F1 (Hao et al., 2011), which measures the number of pages where the attributes are accurately predicted. [^6] Notably, SWDE primarily focuses on identifying textual HTML nodes containing attribute values rather than exact text spans, so we use token-level F1 to identify the most relevant HTML node for each extracted attribute.
4.2 Baselines & Implementation Details
We evaluate the capability of multiple LLMs to perform Schema-Driven IE, including API-based GPT-4 and GPT-3.5 models and open-source models, such as Llama2-Chat-13B (Touvron et al., 2023b), CodeLlama-instruct-13B (Rozière et al., 2023), StarCoder-15.5B (Li et al., 2023), LLaMA-7B (Touvron et al., 2023a), and Alpaca-7B (Taori et al., 2023). We also frame Schema-Driven IE as a TableQA problem, applying multi-choice and extractive QA prompts for template selection and cell attribute prediction, respectively. Furthermore, we also evaluate T5-11B (Raffel et al., 2020) and TaPas (Herzig et al., 2020), a table-specialized LM. For implementation details of InstrucTE and other methods, see Appendix D.
For DisCoMat and SWDE, we compare InstrucTE with established baselines, which either design task-specific architectures, such as FreeDom (Lin et al., 2020) and LANTERN (Zhou et al., 2022), or use LMs pretrained on tables or web pages, like TaPas (Herzig et al., 2020), TaBERT Yin et al. (2020), and MarkupLM (Li et al., 2022).
Figure 4: Ablation studies on various components of our InstrucTE (w/ code-davinci-002) on the ML tables. Interestingly, excluding the table caption improves performance. Our detailed analysis in Appendix E reveals that low-quality captions (e.g., lack of specificity) may confuse the model, leading to inaccurate predictions.
4.3 Main Results
Figure 3 presents the main results from the comparison between InstrucTE and other methods on our Schema-to-Json benchmark. We observe that InstrucTE, in conjunction with API-based models, achieves strong performance across domains and input formats, without any domain-specific labels. With GPT-4, InstrucTE can outperform fine-tuned models on ML and chemistry tables. However, a substantial disparity remains compared to human performance, e.g., the Table-F1 on double-annotated examples for ML tables stands at 96.6 when applying thresholded token-level F1 for attribute matching, which is 22.4 F1 points higher than GPT-4.
For DisCoMat and SWDE, GPT-4 performs on par or slightly trails behind the fully supervised state-of-the-art methods, signifying the potential of LLMs to act as flexible, powerful tools for extracting information from tables across diverse data formats and domains.
Despite a noticeable gap when compared to API-based LLMs, open-source models, like CodeLlama-instruct-13B, show promising results in ML and web domains, achieving 60.0 Table-F1 and 91.7 Page-F1 on ML tables and SWDE, respectively. This success is likely due to these domains being more represented in the model’s pre-training corpus, suggesting that enhancing the pre-training data in less represented domains, such as chemistry and material science, may be an avenue for narrowing the gap with API-based models.
4.4 Ablation Studies
We assess the impact of different components of InstrucTE, including task formulation and error recovery, using ML tables.
LLMs & Task Formulation
In Table 2, we compare different LLMs, leading to two principal observations. First, code models show strong performance on Schema-Driven IE. This is evident from several key comparisons, such as the performance similarity between code-davinci-002 and GPT-4, the superior performance of code-davinci-002 compared to other GPT-3.5 models, and the fact that CodeLlama-instruct-13B significantly outperforms Llama2-chat-13B, approaching the performance of gpt-3.5-turbo. This superiority of code models might be attributed to their alignment with Schema-Driven IE, which involves converting table source code into JSON records. Second, non-code open-source models with similar sizes (for instance, those in the 6-7B range) tend to achieve comparable fine-tuning performance, though they might exhibit variations in prompting performance.
Subsequently, we compare three task formulations: Schema-to-Json, TableQA, and Function Calling, which is a feature provided by the OpenAI API. [^7] https://platform.openai.com/docs/guides/function-calling In Function Calling, the schema is formatted as function definitions with attributes serving as arguments. The LM is then tasked with selecting the function and generating JSON objects for extracted arguments on a cell-by-cell basis. From the T5-11B fine-tuning experiments, we observe that Schema-to-Json attains better performance than TableQA, demonstrating the value of integrating task-specific instructions and extraction schema in the input. Function Calling with gpt-3.5-turbo shows limited effectiveness, and error analysis suggests that this shortfall primarily stems from the model’s struggle in selecting the correct function. [^8] This finding is supported by a marked performance increase to 63.8 Table-F1 when the gold function is pre-specified. As each function call yields only one JSON object, this method requires cell-by-cell prompting, which is cost-intensive with GPT-4. Due to API budget constraints, our experiments are limited to gpt-3.5-turbo.
Exp. Setup Formulation Model Token-F1 EM
Fine-tuning (# Train=1169) TableQA TaPas (large) 27.7 21.6
T5 (11B) 61.2 46.2
Sche2JSON GPT-J (6B) 49.6 38.4
LLaMA (7B) 51.3 38.0
Alpaca (7B) 50.2 39.4
T5 (11B) 64.1 50.2
No Fine-tuning TableQA Flan-T5 (11B) 36.9 27.7
Func. Calling gpt-3.5-turbo (0613) 22.4 18.4
Sche2JSON GPT-J (6B) 18.6 16.2
LLaMA (7B) 13.5 11.5
Alpaca (7B) 26.8 21.1
Llama2-chat (13B) 31.5 23.0
StarCoder (15.5B) 41.2 32.3
CodeLlama-instruct (13B) 60.0 44.0
gpt-3.5-turbo (0613) 64.1 47.9
text-davinci-003 67.4 50.4
code-davinci-002 72.3 57.6
gpt-4 (0613) 74.2 58.1
Table 2: Test set performance on ML tables with different LLMs and task formulations.
Prompt Components & Error Recoverery
Figure 4 shows InstrucTE’s performance subject to the exclusion of varying prompt components. We use code-davinci-002 for these experiments considering API budget limitations and its resemblance to GPT-4 in terms of performance and context length. We observe that removing supplementary text degrades performance. Table headers contribute positively as expected, while captions surprisingly do not. Further analysis on table captions is provided in Appendix E, which suggests that unclear captions can sometimes mislead the model, resulting in inaccurate predictions. Notably, discarding the extraction schema, specifically JSON templates, causes a substantial performance decline, primarily due to attribute name mismatches in the evaluation. Lastly, we show that InstrucTE’s performance drops significantly without error recovery. Compared to cell-by-cell prompting, error recovery offers similar performance at a fraction of the API cost (\$100 v.s. \$670 on Azure). [^9] The pricing for code-davinci-002 on Azure is \$0.1 per 1,000 tokens as of June 23rd, 2023.
4.5 Performance Analysis
To gain more insights about InstrucTE’s performance, we now discuss a set of key questions.
What types of tables can be better processed by InstrucTE?
We analyze InstrucTE’s performance against table complexity on ML tables, considering two factors: 1) the total cell count, and 2) the number of multi-row/column cells, gauging structural complexity. The Pearson correlation coefficients with InstrucTE’s performance (using GPT-4) are 0.096 and 0.128, respectively, suggesting the model’s effective processing across various table sizes and complexities. Qualitative analysis reveals that InstrucTE performs best with tables containing common terminologies and standard content, particularly when contextually supported by the surrounding text. For example, it accurately identifies "Result" cells featuring well-known datasets (e.g., SQuAD) and metrics (e.g., F1 score). Yet, the challenge lies in quantifying the familiarity of terminology and content, given the model’s black-box nature, and addressing it represents a compelling direction for future research.
How does the data format impact InstrucTE’s performance?
The variation in model performance across datasets from different domains with unique formats raises questions about the influence of format differences. To address this, we conducted experiments converting ML tables from  to HTML and chemistry tables from XML to CSV, utilizing both commercial (tableconvert [^10] https://tableconvert.com/api/) and open-source (TeX4ht [^11] https://tug.org/tex4ht/) tools, and selecting the one with the highest conversion accuracy. Despite tableconvert showing superior conversion quality, residual code from the original formats in the converted tables, e.g.,  commands in HTML tables, presents a novel "code-switching" challenge for InstrucTE. Performance evaluation with GPT-4 reveals a minimal drop for ML tables (from 74.2 to 74.1 in Table-F1) and a more significant decrease for chemistry tables (from 83.4 to 78.1 in Table-F1). Both conversion noise and the model’s format-specific processing capabilities could contribute to these differences. The optimal performance on original formats underlines the necessity of developing models adept at handling diverse data formats directly, rather than relying on format conversion tools.
4.6 Knowledge Distillation
Considering the strong performance of API-based models on Schema-Driven IE, we now show that it is possible to use knowledge distillation (Le et al., 2022; Kang et al., 2023) to build a cost-efficient compact model, using ML tables as a demonstration. Specifically, this process first generates synthetic data by performing inference on unlabeled tables using code-davinci-002, followed by fine-tuning a smaller model (e.g., 7B parameters) using the synthetic data. We compile a collection of 979 arXiv ML papers, submitted between 2008 and 2019, yielding 3,434 tables (containing a total of 100K cells). In Table 3, we can see that LLaMA-7B and Alpaca-7B demonstrate similar performance as seen in the fine-tuning results (Table 2). While fine-tuning LLaMA with LoRA (Hu et al., 2022) presents noticeable computational efficiency, full-parameter fine-tuning of T5-11B matches the teacher model’s performance. [^12] As we observe that the T5-11B student model slightly outperforms the teacher model, we conduct a statistical significance test following Berg-Kirkpatrick et al. (2012). With 1000 bootstrap samples, the p-value is 42.3%, suggesting that the performance gap between the T5-11B student model and the teacher model is not statistically significant.
Model (GPU hours) Token-Level F1 EM
P R F1 P R F1
Teacher code-davinci-002 74.1 71.8 72.3 59.4 56.9 57.6
Student LLaMA-7B (50h) 74.1 67.6 69.1 56.8 53.4 54.3
Alpaca-7B (50h) 72.7 64.8 67.5 56.1 50.0 52.0
T5-11B (380h) 75.8 71.4 73.2 60.3 56.7 58.1
Table 3: Experimental results for knowledge distillation on the ML tables. Student models are trained on 3,434 tables labeled by the teacher model. GPU hours refers to the training time ($\times$ number of GPUs) of student models for one epoch.
4.7 Positive Externalities of InstrucTE
To further validate InstrucTE’s practicality, we integrate it with multi-modal models, like GPT4-V, for extracting data from table images. In an initial study with ML tables, it yields a Table-F1 of 70.2, approaching the 74.2 Table-F1 achieved with the original text inputs. Additionally, we explore InstrucTE’s application to the task of Leaderboard Extraction, where it shows competitive performance against leading supervised systems. Due to space constraints, details on these explorations are provided in Appendix F.
5 Related Work
Table Understanding in NLP Research
Recently there have been many research efforts involving tables, particularly, table-to-text generation (Parikh et al., 2020; Wang et al., 2022; Hu et al., 2023). For example, ToTTo Parikh et al. (2020) introduced the task of open-domain table-to-text generation. Another burgeoning area is question answering and fact verification on tables (Jauhar et al., 2016; Zhong et al., 2017; Yu et al., 2018; Schlichtkrull et al., 2021). Various approaches have emerged, such as semantic parsing for compositional question answering (Pasupat and Liang, 2015), and symbolic reasoning for fact verification (Chen et al., 2020). In contrast, our work transforms tables into structured JSON records, where a data schema is the only supervision provided.
Pre-training on Semi-structured Data
The rise of pre-trained language models Devlin et al. (2019), has stimulated interest in pre-training on semi-structured data. TaPas Herzig et al. (2020) and TaBERT Yin et al. (2020) pre-train on linearized tables with a specialized cell index embedding. TABBIE Iida et al. (2021) employs dual transformers for separate row and column encoding. HTLM Aghajanyan et al. (2022) uses an HTML-specialized pre-training objective, facilitating a novel structured prompting scheme. Similar to our work, TabLLM (Hegselmann et al., 2023) uses general-purpose LLMs to process linearized tables, but we focus on schema-driven IE rather than table classification or question answering.
IE from Semi-structured Data
Information extraction from semi-structured data has gained increasing interest (Carlson and Schafer, 2008; Dong et al., 2020; Gupta et al., 2022; Lou et al., 2023). Innovations such as OpenCeres (Lockard et al., 2019) and ZeroShotCeres (Lockard et al., 2020) highlight open-domain extraction from web data, while AxCell (Kardas et al., 2020) and TDMS-IE (Hou et al., 2019) focus on leaderboard extraction from ML tables. DisCoMat (Gupta et al., 2022) showcases material composition extraction from scientific tables. Despite these advancements, most existing methods require supervised datasets or specialized models for task-specific fine-tuning. Our approach stands out by using LLMs to accurately extract data across various formats and domains without relying on labels or custom extraction pipelines.
6 Conclusion
This paper explores the capabilities of LLMs for extracting structured data from heterogeneous tables. We introduce a new task, Schema-Driven Information Extraction, which converts tables into structured records guided by a human-authored data schema. To facilitate this task, we present a benchmark, comprised of tables from four diverse domains, and evaluate various LLMs through our proposed method InstrucTE. The experiments reveal that while API-based models excel across domains and formats, open-source models display significant potential in specific areas. Moreover, we conduct detailed ablation studies and analyses to investigate the factors contributing to model success, and validate the feasibility of building more compact models through distillation to reduce dependency on APIs. These contributions lay a solid foundation for the continued advancement of efficient open-source table data extraction.
Limitations
While InstrucTE showcases strong performance as an instruction-based prompting approach, it encounters specific challenges. Firstly, its effectiveness is highly dependent on how prompts are phrased. Despite offering guidelines for crafting prompts in Appendix A, e.g., emphasizing the use of clear attribute names and the selection of simpler attributes, a systematic method for developing robust extraction schemas for new domains remains elusive, often relying on iterative experimentation. Additionally, the model’s varying performance across different domains and formats is difficult to interpret, possibly due to biases in the pretraining corpus, a factor we cannot fully analyze due to the opaque nature of the pre-training process. InstrucTE also faces difficulties with dataset-specific nuances, as it operates on general task descriptions without detailed examples, making it challenging to navigate boundary cases effectively.
Beyond the model’s inherent limitations, the availability of specific API-based backbones like GPT-4 and code-davinci-002 may change, impacting reliance on these resources. To reduce this dependency, we include results from open-source models and investigate knowledge distillation as a viable alternative, showing promising results. Our benchmark aims to facilitate future research focused on enhancing smaller, openly accessible models, recognizing the importance of such developments for practical application and broader accessibility.
Ethical Considerations
Our use of OpenAI’s API-based models to distill open-source table extractors complies with OpenAI’s terms of service, as we do not “use the output from the Services to develop models that compete with OpenAI”. Regarding licenses of four datasets in our Schema-to-Json benchmark, the arXiv ML tables align with the licenses of their original papers. The PubMed Chemistry tables, sourced from the PMC Open Access Subset, conform to Creative Commons or equivalent licenses. For the other two datasets, we adapt pre-existing datasets released by the NLP research community, abiding by their respective original licenses.
Acknowledgments
We would like to thank Azure’s Accelerate Foundation Models Research Program and OpenAI’s Researcher Access Program for graciously providing access to API-based models, such as GPT-4. This research is supported in part by the NSF (IIS-2052498), ODNI and IARPA via the HIATUS program (2022-22072200004), and the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001119C0108. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This work is approved for Public Release, Distribution Unlimited.
Appendix A InstrucTE
Prompt Formulation
Our proposed prompt consists of four components: 1) “Input Table (w/ supp. text)” includes the table source code paired with supplementary text from the document; 2) “Extraction Schema” defines the JSON formats for extracted records, encompassing the record type, attribute names, and associated data types; 3) “Task-specific Instructions” outline the task execution process, addressing both the extraction process from individual cells and the traversal strategy across cells, such as “left-right, top-down”; 4) “Initial Record” is used to jump-start the prompting process, including the partial record of the first cell.
For “Input Table (w/ supp. text)”, we employ the BM25 algorithm to retrieve the most relevant paragraphs for each table. For “Extraction Schema”, we propose two guidelines for schema design: 1) Attribute names should be specific, which decreases the probability of the model generating incorrect attributes, or hallucinations. For instance, when extracting relevant attributes about a movie from a movie webpage, it’s advisable to use specific terms such as ‘‘movie name’’ or ‘‘director name’’, rather than the generic ‘‘name’’; 2) Attributes should be strategically ordered, placing simpler attributes ahead of more complex ones as errors in preceding attributes can adversely affect the prediction of subsequent ones due to the autoaggressive nature of LMs. The exact InstrucTE prompts used in our experiments are shown in Table 5 and Table 6.
Cell Detector
We develop a rule-based method to identify numeric cells for both the ML and chemistry tables. Specifically, for the ML tables, we use the row separator “\\” and the column separator “&” to divide the table into cells. We then loop over each cell, checking for numeric values after stripping away any stylized text. In cases, where a cell contains multiple numeric values, such as “$0\pm 0$”, we consistently choose the first numeric value. For the chemistry tables, the parsing process is more straightforward, owing to the structured XML format of the table. Here, we iterate over each cell, verifying if it contains a numeric value once stylized text has been removed. The performance of our rule-based cell detector on two datasets is presented in Table 4. In the case of DisCoMat, we use the cell detector provided by the original paper Gupta et al. (2022).
Dataset Split P R F1
ML Tables Dev 100.0 97.0 98.0
Test 99.9 99.6 99.7
Chem. Tables Dev 100.0 100.0 100.0
Test 100.0 98.3 99.2
Table 4: Results of (numeric) cell detection on ML and chemistry tables.
Dataset Full Prompt
ML Tables
[Retrieve paragraphs]
[Input table]
Here are JSON templates for four types of numeric cells: “Other”, “Result”, “Data Stat.”, and “Hyper-parameter/Architecture”:
{“value”: “xx”, “type”: “Other”}
{“value”: “xx”, “type”: “Result”, “task”: “xx”, “metric”: “xx”, “training data/set”: “xx”, “test data/set”: “xx”, “model/method”: “xx”, “model/method settings”: {“xx”: “yy”}, “experimental settings”: {“xx”: “yy”}}
{“value”: “xx”, “type”: “Data Stat.”, “dataset”: “xx”, “attribute name”: “xx”, “sub-set/group name”: “xx”, “dataset features”: {“xx”: “yy”}}
{“value”: “xx”, “type”: “Hyper-parameter/Architecture”, “model”: “xx”, “parameter/architecture name”: “xx”, “dataset”: “xx”}
Please describe all numeric cells in the above latex table following the JSON templates (proceeding by row in a left-right, top-down direction). For each cell, output one JSON description per line. For any unanswerable attributes in the templates, set their value to the placeholder “xx” if it is of string type and {“xx”: “yy”} if it is of dictionary type.
Cell Description:
{“value”: “[Query cell]”, “type”:
Chem. Tables
[Input table]
Here are JSON templates for six types of numeric cells: “Other”, “IC50”, “EC50”, “CC50”, “MIC”, and “GI50”:
{“value”: “xx”, “type”: “Other”}
{“value”: “xx”, “type”: “IC50”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
{“value”: “xx”, “type”: “EC50”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
{“value”: “xx”, “type”: “CC50”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
{“value”: “xx”, “type”: “MIC”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
{“value”: “xx”, “type”: “GI50”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
Please describe all numeric cells in the above XML table following the JSON templates (proceeding by row in a left-right, top-down direction). For each cell, output one JSON description per line. For any unanswerable attributes in the templates, set their value to the placeholder “xx”.
Cell Description:
{“value”: “[Query cell]”, “type”:
Table 5: InstrucTE prompts used for ML and chemistry tables.
Dataset Full Prompt
DisCoMat
[Input table]
Here are JSON templates for two types of numeric cells: “Other” and “Glass_Compound_Amount”:
{“value”: “xx”, “type”: “Other”}
{“value”: “xx”, “type”: “Glass_Compound_Amount”, “constituent compound name”: “xx”, “unit”: “xx”, “glass material/sample name/id/code”: “xx”}
Please describe all numeric cells in the above table following the JSON templates (proceeding by row in a left-right, top-down direction). For each cell, output one JSON description per line. For any unanswerable attributes in the templates, set their value to the placeholder “xx”.
Cell Description:
{“value”: “[Query cell]”, “type”:
SWDE-auto
[Input webpage]
Here is the JSON template for automobile attribute extraction:
{“webpage title”: “xx”, “automobile model (year)”: “xx”, “price”: “xx”, “engine type”: “xx”, “fuel economy”: “xx”}
Please extract the automobile’s attributes from the HTML code above following the JSON template. For any unanswerable attributes in the template, set their value to the placeholder “<NULL>”.
{“webpage title”: “[webpage title]”, “automobile model (year)”:
Table 6: InstrucTE prompts used for DisCoMat and SWDE. For SWDE, we use the “Auto” vertical as an illustrative example, and the prompts for other verticals differ only in attribute names (refer to Table 7 for the attributes of each vertical).
Vertical # Sites # Pages Attributes
Auto 10 17,923 model, price, engine, fuel-economy
Book 10 20,000 title, author, ISBN-13, publisher, publish-date
Camera 10 5,258 model, price, manufacturer
Job 10 20,000 title, company, location, date
Movie 10 20,000 title, director, genre, rating
NBA Player 10 4,405 name, team, height, weight
Restaurant 10 20,000 name, address, phone, cuisine
University 10 16,705 name, phone, website, type
Table 7: SWDE statistics.
Appendix B arXiv Machine Learning Tables
Extraction Attributes
We design a set of extraction attributes for each of the three primary types of numeric cells in ML tables: “Result”, “Hyper-parameter”, and “Data Statistics”. These attributes are outlined in detail below.
  • “Result” includes seven attributes: training data, test data, task, metric, model, model settings and experimental settings. The first five attributes are fixed, with answers being text spans in the paper. The last two attributes, model settings and experimental settings, are free-form attributes, with answers being JSON objects. For example, the experimental settings attribute may be {“number of training examples”: “0”} for a zero-shot setting. This scheme is more detailed than previous approaches (Hou et al., 2019; Kardas et al., 2020) and can accommodate a broader range of ML paradigms and provide more granular information.
  • “Hyper-parameter” includes optimization parameters like learning rate and batch size, as well as numeric descriptions of model architectures such as layer count. The three fixed attributes for this category are: model, parameter/architecture, and dataset.
  • “Data Stat.” covers four attributes: dataset, dataset attribute, sub-set/group, and dataset features. The sub-set/group specifies a dataset subset (e.g., “train” or “test”), while dataset features, a free-form attribute, captures various dataset characteristics like the language or domain.
Annotation Process
We sample 10 papers from each of three pertinent arXiv fields: Machine Learning, Computer Vision, and Natural Language Processing. After removing papers without source code or any tables, a total of 25 papers are covered in our dataset. To optimize the annotation budget and the dataset diversity, we cap the number of annotated tables to five per paper. Recognizing the domain-specific expertise needed, we employ expert annotators with backgrounds in ML research, who are provided with tables in both  and PDF formats and encouraged to thoroughly read the paper before annotation. The annotation process comprises two steps: 1) identifying the numeric cells and their record types, and 2) filling in the slots of pre-determined attributes, forming a JSON record with keys as attribute names and values as extracted content, in a text editor. Consequently, the dataset contains 122 tables, with 3,792 cells and 21K attributes annotated.
Figure 5: Results of comparing various metrics, including token-level F1, SBERT, and BERTScore, to human judgment over different thresholds on ML tables. Numbers are computed over 677 sampled attributes that are paired with respective gold references.
Figure 6: An error analysis of edge cases in which the predictions made by InstrucTE with captions default to “Other” (resulting in an 0 F1). Our hypothesis that this issue may stem from the caption’s lack of specificity is tested by manually expanding the caption (displayed on the right). This amendment significantly improves the performance on these edge cases, increasing the F1 score to 92.3.
Figure 7: An example of Table-F1 calculation, where two predicted records are compared against the two gold records.
Appendix C Evaluation Metrics
Comparing an LLM-predicted JSON object with a gold JSON object is a non-trivial task, as those generative LLMs may produce text spans that do not exactly exist in the input table. Consequently, we devote substantial effort to examining various metrics to determine the one best suited for our task using ML tables. Here, we consider three metrics: the standard token-level F1 to capture the level of lexical overlap between the predicted and gold attributes, and two semantic similarity metrics, SBERT (Reimers and Gurevych, 2019) and BERTScore (Zhang et al., 2020), to identify semantically similar expressions (e.g., # params vs. the number of parameters).
Meta Evaluation
To assess how accurate each metric is compared to human evaluation, we manually annotated predicted-gold attribute pairs as to whether or not each pair matches. We consider a given pair to “match” if they are semantically equivalent, meaning they can be used interchangeably. For attributes that encapsulated multiple sub-attributes, we consider a pair to match if at least half of the sub-attributes are matched (i.e., F1 score $\geq$ 0.5), with the decision for each sub-attribute being based on the same as in the text-span attributes. For the set of pairs to annotate and use as a test set, we sample a total of 100 cell pairs (i.e., 677 attribute pairs) according to the following process: 1) we first uniformly sample a table from the development set (containing 10 papers); and 2) we then sample a random cell from the table, ensuring there were no duplicate cells. For each pair of predicted-gold attributes, each metric’s decision (1 or 0) is made using a specific threshold. For example, if the token-level F1’s score for paired attributes is 0.4 and the threshold is 0.5, then the decision would be 0, indicating no match. The decisions over the test set containing 677 attribute pairs are then compared to human evaluation. In this binary classification problem, F1 is used to evaluate the performance of the metrics.
In Table 8, we present the performances of each metric with the optimal threshold for each. Surprisingly, we find that the token-level F1 (with a threshold of 0.25) decision aligns nearly perfectly with human judgment, and performs the best among all metrics for our task. This might suggest that discerning subtle differences is more crucial than identifying different phrases with the same meaning for this task. Based on these empirical findings, we opt for the token-level F1 for automatic evaluation at the attribute level. This choice is highly desirable not only because of its high accuracy but also due to its simplicity.
token-level F1 SBERT BERTScore
Meta Eval. F1 97.0 95.6 96.7
Threshold 0.25 0.55 0.85
Table 8: Results of comparing various metrics, including token-level F1, SBERT, and BERTScore, to human judgment on ML tables. Numbers are computed over 677 sampled attributes that are paired with gold references. The highest achieved F1 scores are displayed alongside the thresholds. A complete illustration of results, sorted by thresholds, can be found in Figure 5 in Appendix.
Appendix D Implementation Details
Considering the lengthy source code for tables, we employ different strategies to encode the input table and perform Schema-Driven IE, based on the context length of the chosen LLM. For LLMs with a larger context length, such as GPT-4, code-davinci-002, and CodeLlama, we input the full table and conduct the proposed error recovery process. For LLMs with a more limited context length, such as LLaMA and T5-11B, we query each target cell individually. The input table is condensed by rows, retaining the first two rows, typically containing headers, and the row with the query cell, with the token <select> pinpointing the position of the query cell. We use greedy decoding to maximize the reproducibility of our results.
For the TableQA setting, we divide the problem into two steps: selecting the record type and predicting the relevant attributes. For T5 and Flan-T5, the first step is modeled as a multi-choice QA problem, where the model chooses the type of the query cell from a list of provided options. The second step is modeled as an extractive QA task, asking the model to pinpoint the answer spans for the attributes associated with the selected type. For TaPas, the initial step is treated as a classification problem, whereas the latter one is handled as a cell selection problem. The hyper-parameters used for fine-tuning T5 and TaPas are presented in Table 9.
T5 (11B) TaPas
learning rate 1e-4 5e-5
batch size 8 32
# epoches 5 10
Table 9: Hyper-parameters used for fine-tuning T5 and TaPas.
Appendix E Error Analysis of Caption
In Section 4.4, we observe an unexpected finding that table captions do not enhance performance, but rather seem to detract from it, which is counterintuitive. To delve deeper into this observation, we conduct an error analysis. This involves comparing the performances of our InstrucTE system with and without captions at the table level. This analysis uncovers a few outliers (3 out of 68) where including a caption leads to a 0 F1 score, whereas the score is near perfect when the caption is excluded. For instance, as depicted in Figure 6, the predictions all fall into the “Other” category when a caption is included, leading to a 0 F1 score in these outlier instances. Conversely, removing the caption results in an F1 score of 89.3. This high score is due to the fact that retrieved paragraphs provide ample contextual information (e.g., “hate speech detection”) without the presence of a caption.
We hypothesize that the model’s inclination to predict “Other” in the presence of a caption may be a consequence of the captions’ lack of specificity with respect to the attributes relevant to the table cells (for example, “hate speech detection”). This lack of explicit, relevant details could create confusion in associating the caption with the retrieved paragraphs, thereby misleading the model. To test our hypothesis, we manually adjust the captions to include more specific attributes, such as “hate speech detection” and “T5-Base.” As a result, we observe an improvement in the model’s performance with the revised caption, with the total F1 score even exceeding that achieved without a caption. This outcome partially supports our hypothesis and suggests that carefully crafted captions could indeed be beneficial, aligning with our initial expectations. However, this investigation also points to the fact that the model currently lacks robustness in handling these outlier scenarios.
Appendix F Positive Externalities of InstrucTE
F.1 Extraction from Table Images
One practical challenge with InstrucTE is the need for tables in a textual format, while many tables are available only as PDFs or images. To address this, we integrate InstrucTE with multi-modal models to extract structured data from table images. Specifically, we experiment with two strategies: 1) direct extraction from table images, and 2) a pipeline that first employs multi-modal models to transform table images into text, and then run InstrucTE on the textual tables.
In a preliminary study with ML tables, we use GPT-4V as the backbone for InstrucTE. We find that the pipeline method yields a Table-F1 score of 70.2 from image inputs, approaching the 74.2 Table-F1 achieved with the original text inputs. It outperforms direct extraction using GPT-4V, which attains only a Table-F1 score of 46.4, as the pipeline can capitalize on InstrucTE’s error recovery capabilities, resulting in more thorough and accurate extractions.
Additionally, we test IDEFICS-80b-instruct (Laurençon et al., 2023), a leading open-source multi-modal model, which unfortunately could not perform the table-text conversion or direct extraction. [^13] The IDEFICS-80b-instruct model either produces unrelated content or simply output ”I am sorry, but I cannot generate LaTeX code from the table.” This suggests a clear avenue for future research to enhance multi-modal models’ ability to accurately process image-based tables.
Figure 8: Generate  code for image tables using GPT-4V.
F.2 Leaderboard Extraction from ML Papers
Task Definition & SOTA Methods
The task of leaderboard extraction (Hou et al., 2019; Kardas et al., 2020) entails extracting leaderboard tuples (task, dataset, metric, score) from tables in ML papers. Unlike our proposed Schema-Driven IE, which requires open-domain span identification, leaderboard extraction presumes prior knowledge of all leaderboards, represented as pre-defined (task, dataset, metric) tuples, and centers on linking numeric cells to these leaderboards.
The state-of-the-art leaderboard extraction method, AxCell Kardas et al. (2020), is a comprehensive pipeline system comprising four components: Table Type Classification, Table Segmentation, Cell Linking, and Filtering. For each component, except the last one, AxCell employs a supervised model. It starts with table type classification to identify result-related tables, which are then passed to the table segmenter responsible for annotating the header cells of the table. Following this step, a retrieval model links numeric cells in the table to pre-defined leaderboards using human-engineered features. Lastly, AxCell filters and selects the best record based on the leaderboard taxonomy criteria, such as retaining higher values for "Accuracy" and lower ones for "error rate".
Application of InstrucTE
To extract leaderboards from an ML paper, we consider all tables that contain numeric cells, instead of selecting tables via a trained classifier as in AxCell. For each table, we run InstrucTE using a customized leaderboard extraction JSON template. This template resembles the ML-table template with two additional fixed attributes: eval split and eval class in the “Result” cell template. We add the eval split attribute because the evaluated split is essential information for this task; for instance, “dev F1 and “test F1 are treated as different metrics in the leaderboard taxonomy. The eval class attribute is used to exclude sub-set or sub-class results that are typically present in analysis tables. After generating all predicted cell descriptions, we filter them based on three criteria: 1) the type attribute must be “Result”; 2) the eval class attribute must be “all” or “Null” as observed on the development set; and 3) the cell must be bolded in the table, as this usually indicates its superior performance and possible relevance to the leaderboard. For papers without any bolded cells, we experiment with two strategies: 1) include all the remaining cells in the table that meet the first two criteria; 2) use cells selected by AxCell, as its engineered features for cell selection may be useful. This hybrid system is referred to as InstrucTE+. We then use the predicted task, dataset, and metric attributes in each JSON record to match with the pre-defined leaderboards using token-level F1, and we select the leaderboard with the highest average score over three attributes. Finally, following AxCell, we choose the best record based on the leaderboard taxonomy criteria, e.g., retaining higher values for "Accuracy" and lower ones for "error rate".
Method Micro-Average Macro-Average
P R F1 P R F1
AxCell 25.4 18.4 21.3 21.5 21.5 20.0
InstrucTE 20.1 20.8 20.5 20.3 23.1 19.6
InstrucTE+ 23.9 21.2 22.4 21.2 23.7 20.5
Table 10: Leaderboard extraction results on the PWC Leaderboards dataset.
Results
We compare InstrucTE with AxCell on PWC Leaderboards Kardas et al. (2020), the largest dataset for leaderboard extraction. For InstrucTE, we use code-davinci-002 given its excellent performance on Schema-to-Json. Table 10 presents the results of both methods. We can see that InstrucTE achieves competitive performance compared to the supervised AxCell, highlighting the efficacy of our proposed approach. When we enhance InstrucTE with AxCell’s cell selection capabilities to create InstrucTE+, it outperforms AxCell, demonstrating the promising potential of combining these two approaches.